OpenStudy (anonymous):

how do you find the sum of the measures of the interior angles of the polygon

Parth (parthkohli):

Add up all the interior angles :P

OpenStudy (anonymous):

Interior +exterior = 180 n interior = (n-2)180 (n-2) 180 + exterior = 180 exterior are 180(2) = 360

OpenStudy (anonymous):

Alright. So what if my shape has 7 sides how do i find the interior angle?

OpenStudy (anonymous):

the formula is \[2\times(n-2)\times90 degree\]

OpenStudy (anonymous):

as said by @eliassaab .. so you have to put 7 in the place of n

OpenStudy (zepp):